OpenStudy (anonymous):

How high is a tree that casts a 25-ft shadow at the same time a 8-ft post casts a shadow which is 3-ft long?

jimthompson5910 (jim_thompson5910):

hint: solve for x 25/x = 3/8

OpenStudy (anonymous):

66.6

jimthompson5910 (jim_thompson5910):

you are correct the answer is approximately x = 66.67 if you round it to two decimal places

OpenStudy (anonymous):

cross multiply, x/25=8/3 you get 3x=200 divide both sides by 3 and you get x=66.666ft